A compound is composed of 5.045g of carbon, 0.847g of hydrogen and 3.36g of oxygen. Find the empirical formula for this compound knowing that H = 1g/mole , O = 16g/mole and C =12 g/mole.

1
Expert's answer
2022-03-12T02:00:31-0500

"\begin{matrix}
 &C&H&O\\moles&\frac{5.045}{12}&0.847&\frac{3.36}{16}
 \\ratio&2.00&4.03&1.00\\mole\>ratio&2&4&1
\end{matrix}"


Empirical Formula"=C_2H_4O"